OpenStudy (anonymous):

HELP!! I GIVE MEDALS :) Provide an example of a polynomial not in standard form. Explain why it is is not in standard form.

OpenStudy (anonymous):

6+3x^3+7x Because the term with the highest degree is not first.

OpenStudy (anonymous):

3x + 2x^3 - 4x^2 = -12 because you need to rearrange it so it is in standard form of 2x^3-4x^2+3x+12=0 (ax^3-bx^2+cx+d=0)